OpenStudy (kayla0297):

Using exact value symmetry I find that sin330 = +1/2... But shouldn't it be negative?

ganeshie8 (ganeshie8):

good, how did u get 1/2 ?

OpenStudy (kayla0297):

*facepalm* Looking at the wrong question in my book, I mean sqrt3/2.

ganeshie8 (ganeshie8):

yes could you also show me how to get sqrt3/2 ? :)

OpenStudy (kayla0297):

Now you have me worried xD Well here is my working: sin330 = sin(360-30) =+sin30 sin = opp/hyp Using the 30/60 degree triangle, opp=sqrt3 and hyp=2 Which then leads me to believe that I was looking from the 60degree that I got sqrt3/2, but sin30 should be 1/2...

ganeshie8 (ganeshie8):

small mistake

ganeshie8 (ganeshie8):

since the period of `sin` is `360`, adding or subtracting 360 from angle will not change anything, so : sin(330) = sin(360 - 30) = sin(-30) = -sin(30)

ganeshie8 (ganeshie8):